OpenStudy (anonymous):

Can you add the square root of 6 and the square root of 24? Explain why or why not.

OpenStudy (anonymous):

24 can be written as 6 * 4. If we take the root of this we get \[\sqrt{6\times4} = 2\sqrt{6}\] We can now add the square root of 6 as the question asked to give\[\sqrt{6} \times 2\sqrt{6} = 3\sqrt{6}\]. So the answer is yes...

OpenStudy (anonymous):

Sorry the last equation should read:\[\sqrt{6}+2\sqrt{6}=3\sqrt{6}\]